OpenStudy (anonymous):

Algebra 2 Will give medal

OpenStudy (anonymous):

@LynFran one sec while i post it

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

Remember \[\large acos(bx)+d\]\[\large amplitude~=~a\]\[\large period~=~\frac{2\pi}{b}\]\[\large Vertical~change~=~d\] If the maximum is 4 and the minimum is -2, the distance between them is 6. Half of 6 is 3, so the amplitude would be 3. Since the midline is usually 0 for the cos function, with no vertical change, the maximum and minimum would be 3 and -3 respectively. To boost them both up by 1, we'd make the vertical change +1. Finally the period is \(\large \frac{\pi}{2}\) so we can make the formula \[\large \frac{\pi}{2}~=~\frac{2\pi}{b}\] in order to find out what b is.

OpenStudy (lynfran):

to get the phase shift we take (4x+pi)=0 and slove for x so 4x=-pi x=-pi/4 the period i pi in the tan function, so pi/4